# Forecasting When to Forecast: Accelerating Diffusion Models with Confidence-Gated Taylor

### v2

@@ 4.2 Text-to-image model
Quantitative Study. We evaluate our method under two different settings and compare it with prior works, including PAB [43], Teacache [13], and TaylorSeer [15]. As shown in Table 1, our approach achieves the best trade-off between inference efficiency and generation quality in both settings. Under the low acceleration setting, our method achieves the fastest latency of 11.26s, representing a 2.06× speedup, while also attaining the highest Image Reward (1.0398) and SSIM (0.937), with PSNR (29.42)—substantially outperforming other baselines. Importantly, under this acceleration setting, the Image Reward score remains identical to that of the original model, demonstrating that our method preserves generation quality without compromise. In the more aggressive setting, our method further reduces latency to 7.19s, with a 3.13× speedup, maintaining competitive visual quality (Image Reward 0.9956, SSIM 0.813, PSNR 22.00). Compared to TaylorSeer [15], our method achieves a latency reduction of 1.02 seconds, while simultaneously improving SSIM by 25.5% and PSNR by 26.4%.
Qualitative Study. Figure 3 shows that our method preserves high image quality even under accelerated inference. The generated images retain the essential subjects and background elements from the original outputs. In contrast, other methods often yield blurry results or common-sense inconsistencies, such as lamps appearing on rooftops, an incomplete truck in Teacache [13], or a fuzzy bed in Taylorseer [15]. These results highlight our method’s ability to achieve a better trade-off between inference speed and visual fidelity.
@@ 4.3 Class-to-Image model
Quantitative Study. To further validate the generality of our method, we conduct experiments on DiT using various baseline acceleration strategies. As shown in Table 2, our method offers a superior balance between inference speed and visual quality. Specifically, our method attains the fastest inference latency (0.181s) and the highest speed-up ratio (2.36×) while preserving visual quality comparable to the original 50-step generation. In terms of quality metrics, we achieve an FID of 2.34, sFID of 4.28, outperforming most efficient baselines such as FORA [28], ToCa [50], and DuCa [49]. Notably, although TaylorSeer achieves a slightly lower FLOPs count, it suffers from higher latency due to its module-wise prediction overhead. In contrast, our method applies a lightweight last block forecast and a dynamic gating mechanism, resulting in lower runtime overhead without sacrificing fidelity. In addition, our method preserves efficient attention mechanisms [4] throughout, ensuring compatibility with common lightweight DiT backbones.
@@ 4.4 Text-to-Video model
Quantitative Study. Quantitative results on the Wan Video model are shown in Table 3. Our method achieves the highest inference speed, with a latency of only 40.02 seconds, representing a 4.14× acceleration over the original 50-step baseline (165.69s). Notably, this speedup outperforms all other training-free acceleration methods, including Teacache(3.83×) [13] and Taylorseer(3.66×) [15], while also offering superior visual quality. Specifically, our approach maintains a VBench score of 79.01%, just 1.3% below the original unaccelerated baseline (80.31%), yet significantly higher than PAB (71.86%) [43], Teacache (75.39%) [13], and TaylorSeer (76.68%) [15]. In terms of perceptual quality, our method achieves the best SSIM (0.5495) and PSNR (16.04), which outperform TaylorSeer by 63.5% and 31.6% respectively, demonstrating that our adaptive caching mechanism helps reduce approximation errors and better preserve visual details. These results suggest that our method not only enables the highest acceleration but also minimizes quality degradation, effectively balancing speed and fidelity.
@@ 4.5 Ablation Study
Table 3: Quantitative comparison of our method and baselines on the Wan model for text-to-video generation.
Method Acceleration Visual Quality
Wan Video [33] Latency(s)↓ Speed↑ VBench(%)↑ SSIM ↑ PSNR↑ LPIPS ↓
ori 50steps 165.69 1.00x 80.31 - - -
PAB [43] 78.02 2.12x 71.86 0.4162 15.9141 0.5074
Teacache ($\delta=0.26$[13] 43.25 3.83x 75.39 0.3398 12.1620 0.6275
Taylorseer ($N=5$, O=1O=1[15] 45.27 3.66x 76.68 0.3338 12.1926 0.5902
Ours ($\epsilon=0.13$) 40.02 4.14x 79.01 0.5495 16.0418 0.3185
Performance of Different Prediction Confidence Threshold $\epsilon$: Figure 1 plots the speed–quality trade‑off of our method as the confidence threshold $\epsilon$ varies ($\epsilon\in{0.13,0.30,0.50,0.80}$). As expected, larger thresholds admit more aggressive reuse and thus yield higher speed‑up ratios. Although image‑quality metrics (VBench, SSIM, PSNR) gradually decline with increasing $\epsilon$, our curves remain above those of all baselines at comparable speeds, indicating a slower quality‑decay rate. The advantage is most pronounced relative to PAB: once PAB approaches a 2× speed‑up, its quality drops sharply, whereas our method still delivers substantially higher fidelity. These results confirm that the Prediction Confidence Gating mechanism provides a flexible knob for trading speed for quality while preserving a consistently better balance than competing approaches.
Table 4: Ablation study of each component in our method on the FLUX model for text-to-image generation.
Forecast level PCG Latency(s)↓ PSNR↑ Image Reward↑ SSIM ↑ Memory Usage (GB)↓
module 13.56 24.02 1.0005 0.850 39.69
last block 12.41 24.00 1.0109 0.849 35.73
last block 11.26 29.42 1.0398 0.937 35.78
@@
Figure 4: Correlation between the first block and the last block Taylor errors.

### v3

@@ 4.2 Text-to-image model
Quantitative Study. We evaluate our method under two different settings and compare it with prior works, including PAB [zhao2024real], Teacache [liu2025timestep], and TaylorSeer [liu2025reusing]. As shown in Table 1, our approach achieves the best trade-off between inference efficiency and generation quality in both settings. Under the low acceleration setting, our method achieves the fastest latency of 11.26s, representing a 2.06× speedup, while also attaining the highest Image Reward (1.0398) and SSIM (0.937), with PSNR (29.42)—substantially outperforming other baselines. Importantly, under this acceleration setting, the Image Reward score remains identical to that of the original model, demonstrating that our method preserves generation quality without compromise. In the more aggressive setting, our method further reduces latency to 7.19s, with a 3.13× speedup, maintaining competitive visual quality (Image Reward 0.9956, SSIM 0.813, PSNR 22.00). Compared to TaylorSeer [liu2025reusing], our method achieves a latency reduction of 1.02 seconds, while simultaneously improving SSIM by 25.5% and PSNR by 26.4%.
Furthermore, we evaluate the GenEval benchmark on the FLUX model. The results are shown in Table 2. It demonstrates that different methods achieve relatively close performance on GenEval. However, our approach consistently demonstrates strong competitiveness under both low and high acceleration ratios. This further validates the effectiveness of our method.
Qualitative Study. Figure 4 shows that our method preserves high image quality even under accelerated inference. The generated images retain the essential subjects and background elements from the original outputs. In contrast, other methods often yield blurry results or common-sense inconsistencies, such as lamps appearing on rooftops, an incomplete truck in Teacache [liu2025timestep], or a fuzzy bed in Taylorseer [liu2025reusing]. These results highlight our method’s ability to achieve a better trade-off between inference speed and visual fidelity.
@@ 4.3 Class-to-Image model
Quantitative Study. To further validate the generality of our method, we conduct experiments on DiT using various baseline acceleration strategies. As shown in Table 3, our method offers a superior balance between inference speed and visual quality. Specifically, our method attains the fastest inference latency (0.181s) and the highest speed-up ratio (2.36×) while preserving visual quality comparable to the original 50-step generation. In terms of quality metrics, we achieve an FID of 2.34, sFID of 4.28, outperforming most efficient baselines such as FORA [selvaraju2024fora], ToCa [zou2024accelerating], and DuCa [zou2024acceleratingduca]. Notably, although TaylorSeer achieves a slightly lower FLOPs count, it suffers from higher latency due to its module-wise prediction overhead. In contrast, our method applies a lightweight last block forecast and a dynamic gating mechanism, resulting in lower runtime overhead without sacrificing fidelity. In addition, our method preserves efficient attention mechanisms [dao2022flashattention] throughout, ensuring compatibility with common lightweight DiT backbones.
Table 6: Performance comparison under different prediction confidence thresholds $\epsilon$.
Confidence Thresholds $\epsilon$ Latency(s)↓ Speed↑ Image Reward↑ SSIM↑ PSNR↑
0.03 11.26 2.06x 1.0398 0.937 29.42
0.08 7.95 2.92x 1.0220 0.862 24.22
0.13 7.19 3.23x 1.0216 0.813 22.00
0.18 6.13 3.79x 0.9994 0.773 20.81
@@ 4.4 Text-to-Video model
Quantitative Study. Quantitative results on the Wan Video model are shown in Table 4. Our method achieves the highest inference speed, with a latency of only 40.02 seconds, representing a 4.14× acceleration over the original 50-step baseline (165.69s). Notably, this speedup outperforms all other training-free acceleration methods, including Teacache (3.83×) [liu2025timestep] and Taylorseer (3.66×) [liu2025reusing], while also offering superior visual quality. Specifically, our approach maintains a VBench score of 81.86%, just 0.9% below the original unaccelerated baseline (82.72%), yet significantly higher than PAB (73.60%) [zhao2024real], Teacache (76.67%) [liu2025timestep], and TaylorSeer (78.71%) [liu2025reusing]. In terms of perceptual quality, our method achieves the best SSIM (0.6319), PSNR (17.2108), and LPIPS (0.2882), which outperform TaylorSeer by 34.06%, 26.36%, and 49.94%, respectively, demonstrating that our adaptive caching mechanism helps reduce approximation errors and better preserve visual details. These results suggest that our method not only enables the highest acceleration but also minimizes quality degradation, effectively balancing speed and fidelity.
Quantitative results on the Hunyuan Video model are shown in Table 5. Consistent with the results on Wan Video, our method achieves the fastest inference speed on Hunyuan Video while maintaining high visual quality. Specifically, it reaches a 3.70× speedup with an inference latency of 27.45s, outperforming baselines such as TaylorSeer, TeaCache, and PAB. The VBench score drops minimally compared to the original model, and perceptual metrics (SSIM, PSNR, LPIPS) indicate that our approach also provides superior visual fidelity, with significant improvements over TaylorSeer. Overall, experiments on these two text-to-video models further validate the effectiveness and generalizability of our method.
@@ 4.5 Ablation Study
Figure 5: Number of denoising steps below and above the $\epsilon$ threshold for different $\epsilon$ values.
Performance of Different Prediction Confidence Thresholds $\epsilon$: Table 6 plots the speed–quality trade‑off of our method as the confidence threshold $\epsilon$ varies ($\epsilon\in{\color[rgb]{0,0,0}\{0.03,0.08,0.13,0.18\}}$). As expected, larger thresholds admit more aggressive reuse and thus yield higher speed‑up ratios, while a smaller $\epsilon$ results in slower inference but better preservation of image quality. This controllable trade-off allows us to select manually a balanced $\epsilon$ that achieves both substantial acceleration and high output fidelity. Furthermore, to better illustrate how $\epsilon$ influences acceleration, we report Figure 5 showing the number of steps exceeding or falling below the $\epsilon$ threshold at different values. As shown in the figure, with a larger $\epsilon$, more steps fall below the threshold, which means that during inference, a greater portion of steps rely on Taylor-predicted features of the last block. Consequently, the acceleration effect becomes more pronounced. Although image‑quality metrics (Image Reward, SSIM, PSNR) gradually decline with increasing $\epsilon$, our curves remain above those of all baselines at comparable speeds, indicating a slower quality‑decay rate as is shown in Figure 1. The advantage is most pronounced relative to PAB: once PAB approaches a 2× speed‑up, its quality drops sharply, whereas our method still delivers substantially higher fidelity. These results confirm that the Prediction Confidence Gating mechanism provides a flexible knob for trading speed for quality while preserving a consistently better balance than competing approaches.
Table 7: Performance comparison between constant and timestep-dependent thresholds $\epsilon$.
Confidence Thresholds $\epsilon$ Latency(s)↓ Speed↑ Image Reward↑ SSIM↑ PSNR↑
0.08 7.95 2.92x 1.0220 0.862 24.22
adaptive $\epsilon$ 8.11 2.86x 1.0339 0.904 26.91
Table 8: Ablation study of each component in our method on FLUX for text-to-image generation.
Forecast level PCG Latency(s)↓ Image Reward↑ PSNR↑ SSIM ↑ Memory Usage (GB)↓
module 13.56 1.0005 24.02 0.850 39.69
last block 12.41 1.0109 24.00 0.849 35.73
last block 11.26 1.0398 29.42 0.937 35.78
Table 9: Performance comparison when different blocks are used as prediction indicators.
Prediction Indicator Latency(s)↓ Speed↑ Image Reward↑ SSIM↑ PSNR↑
First Block 7.19 2.92x 1.0217 0.813 22.00
Third Block 7.49 2.86x 1.0170 0.801 21.77
In addition, motivated by [wu2024diffusion], we also experimented with a timestep-dependent $\epsilon$. $\epsilon$ is set smaller at early timesteps and larger at later ones, since early denoising determines global structure while later steps mainly refine details. As reported in Table  7, the timestep-dependent $\epsilon$ schedule achieves better trade-offs. At nearly identical inference speeds (only 0.16s difference), it improves image reward, SSIM, and PSNR, with PSNR increased by 11.2%. This demonstrates that our framework is flexible and can naturally benefit from adaptive thresholding strategies. While our main method adopts a constant $\epsilon$ for simplicity and clarity, we believe that designing principled adaptive schemes is a promising future direction.
@@
Figure 6: Correlation between the first block and the last block of Taylor errors.
Choice of Prediction Indicator. In our experiments, we use the first block as the indicator to assess the accuracy of Taylor prediction. The rationale is that the first block provides a lightweight signal, thereby minimizing the overhead for confidence evaluation. To validate this design choice, we further conduct an ablation study by selecting the third block as the prediction indicator. The results, reported in Table 9, show that the visual quality remains nearly unchanged compared to using the first block, with SSIM differing by only 0.13. However, the inference speed decreases by 4.17%. This confirms that using the first block strikes the optimal balance—it serves as an effective predictor while incurring less runtime cost.